OpenStudy (anonymous):

What is the center and radius of the following circle: (x – 5)2 + (y + 2)2 = 16

OpenStudy (anonymous):

(x-a)^2 +(y-b)^2 = r^2 (a,b) is the center r = radius

OpenStudy (mertsj):